[ If Eric runs 3/4 of a mile every 8 minutes, how much does he run in one hour? What is the unit rate for this problem? ]
Answer by TimothyLamb(4379)

r = d / t
r = (3/4 mile)/(8 min) * (60 min/hr) = 5.625 mph
---
d = r * t
d = 5.625 mile/hr * 1 hr = 5.625 mile
---
answer:
eric runs 5.625 miles in 1 hours
the unit rate is 5.625 miles/hour
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php